If a card is drawn from a deck, find the probability of getting a red card and a 7. (round to 4 decimal places.)
Answer #1

There are 26 red card and four 7 in a deck of 52 cards.

And there are two red cards which are 7.

We want to find, the probability of getting a red card and a 7.

P(red and 7) = 2/52 = 1/26 = 0.0385

Therefore, required probability is 0.0385